OpenStudy (rissyroo118):

Question 9.9. Solve the equation. |x + 7| = 12 {–19, –5} {–19, 5} {–5, 19} {5, 19}

OpenStudy (anonymous):

|x+7|=12 can be written as x+7=12 and x=5 and -x-7=12 as -x=19 {-19,5}